Two boats 1200 feet apart measures an angle of elevation of 35 degree and 42 degree to the top of a lighthouse. how tall is the lighthouse?


1
Expert's answer
2022-05-24T23:19:57-0400

Let "h=" the height of lighthouse.

The ditance from the first boat to the foot of the lighthouse is "\dfrac{h}{\tan 35\degree}."


The distance from the second boat to the foot of the lighthouse is "\dfrac{h}{\tan 42\degree}."

Then


"\dfrac{h}{\tan 35\degree}-\dfrac{h}{\tan 42\degree}=1200"


"h=\dfrac{1200\tan 35\degree\tan 42\degree}{\tan 42\degree-\tan 35\degree}=3779(ft)"

"3779ft=1152m"

I think the answer has no sense.